OpenStudy (anonymous):

What are the solutions of 6x2 - x - 2 = 0? Answer x = 1 over 2 , x = -2 over 3 x = -1 over 2 , x = 2 over 3 x = 1 over 6 , x = -1 x = -1 over 6 , x = 2

OpenStudy (campbell_st):

factorise the quadratic (2x +1)(3x-2) = 0 then x = -1/2 and 2/3

OpenStudy (ash2326):

We have \[6x^2-x-2=0\] A standard quadratic equation is of the form \[ax^2+bx+c=0\] The roots are given by quadratic formula \[x=\frac{-b\pm \sqrt {b^2-4ac}}{2a}\] here we have a=6, b=-1 and c=-2. Let's substitute the values in quadratic formula to find the roots \[x=\frac{-(-1)\pm \sqrt {(-1)^2-4\times 6\times (-2)}}{2\times 6}\] We get \[x=\frac{1\pm \sqrt{1+48}}{12}\] we get \[x=\frac{1\pm \sqrt {49}}{12}=\frac{1\pm 7}{12}\] we get \[x= \frac{1+ 7}{12}, \frac{1-7}{12}\] or \[x= \frac{8}{12}, -\frac{6}{12}\] or \[x=\frac{2}{3}, \frac{1}{2}\]

OpenStudy (ash2326):

Sorry made a typo :( it'd be \[x=\frac{2}{3}, -\frac{1}{2}\]

OpenStudy (anonymous):

lol i was judt about to ask . lol